OpenStudy (kevon18):

Find the function from the given zeros. -1, 2, 1-i

OpenStudy (sooobored):

if a zero is given as "n" the equation equivalent is (x-n)=0

OpenStudy (sooobored):

if we have 3 zeros given as "a" "b" "c" then the equation equivalent is (x-a)(x-b)(x-c)=0

OpenStudy (dumbcow):

\[x = 1-i\] \[x-1 = -i\] \[(x-1)^2 = (-i)^2 = -1\] \[x^2 -2x+1 = -1\] \[x^2 -2x +2 = 0\] \[\rightarrow f(x) = (x+1)(x-2)(x^2 -2x+2)\]